# Parameterized Controls and Isomorphic Representations of G-Expectations

### v1

@@ 4.2 G-SDE Solution via Parameterization
  3. 3.
    The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation
@@
This framework provides a rigorous bridge from classical SDE theory to G-SDEs, leveraging the bounded control structure to transfer solution properties.
@@
Theorem 4.4 (Properties of Solutions ).
The G-SDE solution satisfies:
  1. 1.
    Moment estimates: $\hat{\mathbb{E}}\left[\sup_{0\leq s\leq t}|X_{s}|^{p}\right]<\infty$ for $p\geq 2$
  2. 2.
    Continuous dependence: If $b_{n}\to b$, $\sigma_{n}\to\sigma$ uniformly, then $X_{t}^{n}\to X_{t}$ in $L_{G}^{2}$
4.3 Geometric Brownian Motion — A Clear Example
@@
Through the isomorphism $\Phi$, we transform the G-SDE into a family of classical SDEs:
@@
Note that the two are identical under isomorphism.

### v7

@@ 2. The boundedness of $\Theta$ ensures uniform estimates:
3. The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation ∎
@@
Remark 4.1.
For comparison theorems of G-SDEs, we can directly draw on well-established results from classical SDE theory through our parametric representation. Similarly, a priori estimates and stability analysis for G-SDEs can be obtained by applying the corresponding classical results to each parameterized SDE in the family, provided we maintain uniform estimates with respect to the control parameters. This approach leverages the extensive existing literature on classical stochastic differential equations.
Indeed, numerous fundamental properties and results for classical SDEs—including comparison theorems, a priori estimates, and stability analysis—can be systematically transferred to the G-SDE framework through our parametric representation. The established estimates in classical SDE theory demonstrate consistent behavior across control parameters, making them well-suited for application within our unified framework.
4.2 Geometric Brownian Motion
@@
[…]
@@
Note that the two are identical under homomorphism.